THERMOLUMINESCENCE OF IMPURITY-HELIUM SOLIDS IMMERSED IN LIQUID HELIUM

\underline{R. E. BOLTNEV}, E. B. GORDON, V. V. KHMELENKO, I. N. KRUSHINSKAYA, M. V. MAR-TYNENKO, A. A. PELMENEV, E. A. POPOV, Institute of Energy Problems of Chemical Physics, 142432, Chernogolovka, Moscow Region, Russian Federation; A. F. SHESTAKOV, Institute of Problems of Chemical Physics, 142432, Chernogolovka, Moscow Region, Russian Federation.

A solidification of liquid helium around impurity particles injected in its volume results in IHSP (Impurity-Helium Solid Phase) formation ^{a}, so these particles can be stabilized with inert environment during long time. Particular interest is stabilization of metastable particles. The capture of N(^{2}D) atoms from discharge allows to save ones at T=1.5 K during 10^{4} s - time comparable with their lifetime, 4.4 × 10^{4} s. Nevertheless, even small temperature increase (< 0.1 K) causes the luminescence on the ^{2}D - ^{4}S transition. It was explained as thermoactivated association of neighbouring centers of IHSP, N+N_{2} or N+Rg, which partially removes the prohibition because of ”heavy particle effect” ^{b}. The main part of stabilized nitrogen atoms is in the ground state ^{4}S, so thermoactivated mobility leads to the pair recombination N+N or N+O (O_{2} presents as trace in condensed gas mixture). The results presented in report confirm that heating of IHSP samples causes a blue emission which can be assigned to excited states of N_{2} or NO. Acknowledgement: The work was carried out with support from Russian Foundation for Basic Research (Grant 99-03-33261).

aE. B. Gordon, V. V. Khmelenko, A. A. Pelmenev, E. A. Popov, O. F. Pugachev, A. F. Shestakov, Chem. Phys. 170, 411 (1993).

bR. E. Boltnev, E. B. Gordon, V. V. Khmelenko, I. N. Krushinskaya, M. V. Martynenko, A. A. Pelmenev, E. A. Popov, A. F. Shestakov, Chem. Phys. 189, 367 (1994).